Previous

Calendar

June

2017

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

28

JD148

29

JD149
CU Holiday (Memorial Day)

30

JD150
GLEES Site Visit (JF): Details , Photos (2)

both quad-disk ports are still co-located side-by-side at 22.65m

Yellow Box (SN 950329) connected to clear quad-disk: changed from tilt=20 deg to tilt=40 deg

Gray Box (SN 9301150) connected to Qualmetrics (opaque) quad-disk: tilt=0 deg (no change)

31

JD151

1

JD152

2

JD153

3

JD154

4

JD155

5

JD156
Site Visit (SpB): Details , Photos (30)

Span cylinder at 870 psi. N2 at 750 psi

from XX:XX-XX:XX MST, swiped krypton hygrometer

Problematic thermocouples: T3 (still dead), T4 (still dead)

SpB Jeep

Clear in AM, cloudy in PM

6

JD157
SpB Off

7

JD158
AmeriFlux/Fluxnet workshop (UCB)

8

JD159
AmeriFlux/Fluxnet workshop (UCB)

9

JD160
AmeriFlux/Fluxnet workshop (UCB)

(Full "Strawberry" Moon)

10

JD161

11

JD162

12

JD163

13

JD164
GLEES Site Visit (JF): Details ,

both quad-disk ports are still co-located side-by-side at 22.65m

Yellow Box (SN 950329) connected to clear quad-disk: changed from tilt=40 deg to tilt=10 deg (12.3 deg)

Gray Box (SN 9301150) connected to Qualmetrics (opaque) quad-disk: tilt=0 deg (no change)

SpB Returns

14

JD165
Site Visit (SpB, DB, MG, DE, CF, KG, TM): Details , Photos (20)

Span cylinder at 760 psi. N2 at 600 psi

from XX:XX-XX:XX, zeroed LI-6262 as much as possible

Turned on fans in 6262/6251 boxes

Shutdown snowtemp/noah data loggers

in PM: moved canopy-access towers with DB and crew; Time-Lapse Video (CF)

Problematic thermocouples: T3 (still dead), T4 (still dead)

SpB Jeep

Warm, sunny, a bit windy (snow=gone!)

Flag Day

15

JD166
Site Visit (DB, MG, DE, CF, KG, TM): Details

Installed spectrometer telescope at top of tower (TM, KG, CF)

Continued tower destruction/construction (DB, DE, MG, TM)

16

JD167
Site Visit (SpB, DB, MG, DE, CF, KG, TM): Details , Photos (16)

Group Photo (KG)

changed output of MARK cr23x logger from 10-hz to 1-hz (started mark_12.csi)

Fixed problematic/broken NCTC thermocouples T3 and T4 (borrowed from MARK logger)

DB and crew were building towers; CF and crew were finalizing spectrometer setup

SpB Jeep

Cool in AM, then warm and partly-cloudy

17

JD168

18

JD169
Fathers Day

19

JD170

20

JD171

21

JD172
Summer Solstice

22

JD173

23

JD174

24

JD175

25

JD176

26

JD177

27

JD178

28

JD179

29

JD180

30

JD181

1

JD182